# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v2

@@ Acknowledgement
In Section A, we provide the potential limitations of this work. In Section B, we discuss the societal impacts of our work. In Section C, we introduce some definitions that will be used in the proof. In Section D, we introduce some auxiliary lemma from previous work that we need. In Section E, Section F, Section G, Section H, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version. Finally, in Section I, we provide more experimental results and implementation details.
@@ Appendix B Societal Impact
Appendix C More Notations and Definitions
@@
Definition C.1 ($L_{b}$ (vector) norm).
@@
Definition C.2 ($L_{a,b}$ (matrix) norm).
@@
Appendix D Tools from Previous Work
Section D.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section D.2 introduces the maximum-margin for multi-class.
D.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma D.1 (Lemma 5 in page 8 in [64]).
@@ D.1 Tools from Previous Work: Implying Single/Combined Neurons
D.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma D.2 (Lemma 6 in page 8 in [64]).
@@ D.2 Tools from Previous Work: Maximum Margin for Multi-Class
Appendix E Class-weighted Max-margin Solution of Single Neuron
Section E.1 introduces some definitions. Section E.2 shows how we transfer the problem to discrete Fourier space. Section E.3 proposes the weighted margin of the single neuron. Section E.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section E.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
E.1 Definitions
Definition E.1.
@@ E.1 Definitions
Definition E.2.
Let $\eta$ be defined in Definition E.1. When $k=3$, provided the following conditions are met
@@
E.2 Transfer to Discrete Fourier Space
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@ E.3 Get Solution Set
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
By Lemma E.3, we only need to maximize Equation (E.2).
@@
Consider the odd $p$, Equation (E.2) becomes:
@@
where the first step comes from definition (E.2), the second step follows from Eq. (8), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
E.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition E.5.
@@ E.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition E.6.
Let $\eta$ be defined in Definition E.5. Provided the following conditions are met
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.6.
@@
E.5 Get Solution Set for General $k$ Version
Lemma E.8 (Formal version of Lemma 4.2).
@@ E.5 Get Solution Set for General $k$ Version
  • Let $\Omega_{q}^{{}^{\prime}*}$ be defined as Definition E.6.
@@
By Lemma E.7, we only need to maximize Equation (E.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (E.4) becomes:
@@
where the first step follows from definition (E.4), the second step comes from Eq. (17), the last step follows from Eq. (18), i.e., Euler’s formula.
@@
Appendix F Construct Max Margin Solution
Section F.1 proposed the sum-to-product identities for $k$ inputs. Section F.2 shows how we construct $\theta^{*}$ when $k=3$. Section F.3 gives the constructions for $\theta^{*}$ for general $k$ version.
F.1 Sum-to-product Identities
Lemma F.1 (Sum-to-product Identities).
@@ F.1 Sum-to-product Identities
F.2 Constructions for $\theta^{*}$
Lemma F.2.
@@ F.2 Constructions for $\theta^{*}$
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
Referencing Lemma E.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (F.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
F.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma F.3 (Formal version of Lemma 4.3).
@@ F.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition E.2.
@@
By Lemma E.4, we can get elements of $\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
Appendix G Check Fourier Frequencies
Section G.1 proves all frequencies are used. Section G.2 proves all frequencies are used for general $k$ version.
G.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma E.4,
@@
Setting aside the previously discussed points, it’s established in Lemma D.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
G.2 All Frequencies are Used for General $k$ Version
@@
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma E.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma D.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix H Main Result
Section H.1 proves the main result for $k=3$. Section H.2 proves the general $k$ version of our main result.
H.1 Main result for $k=3$
Theorem H.1.
@@ H.1 Main result for $k=3$
By Lemma E.4, we get the single neuron class-weighted margin solution set $\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.8 and $\gamma^{*}$.
By Lemma F.2 and Lemma D.1, we can construct network $\theta^{*}$ which uses neurons in $\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.8 and Definition 3.5 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma D.2, we know it is the maximum-margin solution.
By Lemma G.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
H.2 Main Result for General $k$ Version
Theorem H.2 (Formal version of Theorem 4.1).
@@ H.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem H.1 by Lemma E.8, Condition 3.8, Lemma F.3, Lemma D.1, Definition 3.5, Lemma D.2, Lemma G.2.
Appendix I More Empirical Details and Results
I.1 Implement Details
@@
I.2 One-hidden Layer Neural Network

### v3

@@ Acknowledgement
In Section A, we provide the potential limitations of this work. In Section B, we discuss the societal impacts of our work. In Section C, we provide more related works. In Section D, we introduce some definitions that will be used in the proof. In Section E, we introduce some auxiliary lemma from previous work that we need. In Section F, Section G, Section H, Section I, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version. Finally, in Section J, we provide more experimental results and implementation details.
@@ Appendix B Societal Impact
Appendix C More Related Work
Theoretical Work About Fourier Transform. To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [37, 36, 43, 44]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [10, 35, 65]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [79, 39, 32, 56, 18, 81, 16, 82, 19, 88].
Appendix D More Notations and Definitions
@@
Definition D.1 ($L_{b}$ (vector) norm).
@@
Definition D.2 ($L_{a,b}$ (matrix) norm).
@@
Appendix E Tools from Previous Work
Section E.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section E.2 introduces the maximum-margin for multi-class.
E.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma E.1 (Lemma 5 in page 8 in [58]).
@@ E.1 Tools from Previous Work: Implying Single/Combined Neurons
E.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma E.2 (Lemma 6 in page 8 in [58]).
@@ E.2 Tools from Previous Work: Maximum Margin for Multi-Class
Appendix F Class-weighted Max-margin Solution of Single Neuron
Section F.1 introduces some definitions. Section F.2 shows how we transfer the problem to discrete Fourier space. Section F.3 proposes the weighted margin of the single neuron. Section F.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section F.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
F.1 Definitions
Definition F.1.
@@ F.1 Definitions
Definition F.2.
Let $\eta$ be defined in Definition F.1. When $k=3$, provided the following conditions are met
@@
F.2 Transfer to Discrete Fourier Space
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@ F.3 Get Solution Set
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
By Lemma F.3, we only need to maximize Equation (F.2).
@@
Consider the odd $p$, Equation (F.2) becomes:
@@
where the first step comes from definition (F.2), the second step follows from Eq. (8), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
F.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition F.5.
@@ F.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition F.6.
Let $\eta$ be defined in Definition F.5. Provided the following conditions are met
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.6.
@@
F.5 Get Solution Set for General $k$ Version
Lemma F.8 (Formal version of Lemma 4.2).
@@ F.5 Get Solution Set for General $k$ Version
  • Let $\Omega_{q}^{{}^{\prime}*}$ be defined as Definition F.6.
@@
By Lemma F.7, we only need to maximize Equation (F.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (F.4) becomes:
@@
where the first step follows from definition (F.4), the second step comes from Eq. (17), the last step follows from Eq. (18), i.e., Euler’s formula.
@@
Appendix G Construct Max Margin Solution
Section G.1 proposed the sum-to-product identities for $k$ inputs. Section G.2 shows how we construct $\theta^{*}$ when $k=3$. Section G.3 gives the constructions for $\theta^{*}$ for general $k$ version.
G.1 Sum-to-product Identities
Lemma G.1 (Sum-to-product Identities).
@@ G.1 Sum-to-product Identities
G.2 Constructions for $\theta^{*}$
Lemma G.2.
@@ G.2 Constructions for $\theta^{*}$
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
Referencing Lemma F.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (G.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
G.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma G.3 (Formal version of Lemma 4.3).
@@ G.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
By Lemma F.4, we can get elements of $\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
Appendix H Check Fourier Frequencies
Section H.1 proves all frequencies are used. Section H.2 proves all frequencies are used for general $k$ version.
H.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma F.4,
@@
Setting aside the previously discussed points, it’s established in Lemma E.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
H.2 All Frequencies are Used for General $k$ Version
@@
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma F.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma E.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix I Main Result
Section I.1 proves the main result for $k=3$. Section I.2 proves the general $k$ version of our main result.
I.1 Main result for $k=3$
Theorem I.1.
@@ I.1 Main result for $k=3$
By Lemma F.4, we get the single neuron class-weighted margin solution set $\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.8 and $\gamma^{*}$.
By Lemma G.2 and Lemma E.1, we can construct network $\theta^{*}$ which uses neurons in $\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.8 and Definition 3.5 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma E.2, we know it is the maximum-margin solution.
By Lemma H.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
I.2 Main Result for General $k$ Version
Theorem I.2 (Formal version of Theorem 4.1).
@@ I.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem I.1 by Lemma F.8, Condition 3.8, Lemma G.3, Lemma E.1, Definition 3.5, Lemma E.2, Lemma H.2.
Appendix J More Empirical Details and Results
J.1 Implement Details
@@
J.2 One-hidden Layer Neural Network